OpenStudy (anonymous):

X, Y, Z and W are integers. The expression X-Y-Z is even and the expression Y-Z-W is odd. If X is even what must be true? W must be odd Y-Z must be odd Z must be odd

OpenStudy (anonymous):

Try it........ u will get that the pattern will repeat.

OpenStudy (anonymous):

Calculate up to A(8) u will see the result.

OpenStudy (anonymous):

i just want to conform its answer ,,,actually my answer is 0 is that correct

OpenStudy (anonymous):

S(1000) = S( 1000 MOD 6)

OpenStudy (anonymous):

what mod 6 does here

OpenStudy (anonymous):

S(1000) = S(4) = 1+2+1-1=3

OpenStudy (anonymous):

Did u see the pattern?

OpenStudy (anonymous):

1,1,0,0,-1,-1 like that

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Nope: S(1)=1 S(2)=3 S(3)=4 S(4)=3 S(5)=1 S(6)=0 S(7)=1 S(8)=3 S(9)=4 S(10)=3

OpenStudy (anonymous):

it was 1 1 0 -1 -1 0 1 1 -1 -1

OpenStudy (anonymous):

The pattern repeats after 6

OpenStudy (anonymous):

you r confusing me

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

First, list A(1), A(2)...A(8)

OpenStudy (anonymous):

Can u do it?

OpenStudy (anonymous):

please give me an answer

OpenStudy (anonymous):